OpenStudy (anonymous):

solve by substitution 3x-2y=5 4x-5y=2

sam (.sam.):

3x-2y=5 4x-5y=2 -------------------------------------------- 3x-2y=5 x=(5+2y)/3 --------------------------------------------- 4x-5y=2. 4(5+2y)/3-5y=2 (4(5+2y))/(3)-5y=2 (4(2y+5))/(3)-5y=2 (4(2y+5))/(3)-5y*(3)/(3)=2 (4(2y+5))/(3)-(15y)/(3)=2 (4(2y+5)-(15y))/(3)=2 (-7y+20)/(3)=2 -7y+20=6 -7y=-14 y=2 ------------------------------------------ x=(5+2y)/3 x=(5+2(2))/3 x=3

OpenStudy (anonymous):

pick either equation and solve for either x or y then use the expression created (for example x=blah) to substitute in the other equation) x+2y=0, will become blah+2y=0 to get the first expression created for x=blah you will eliminate every x, since the only x you can get is the first one in x=... hence you will be left with an equation containing only y and you can solve for y